# Decoherence of Josephson coupling and thermal quenching of the Josephson diode effect in bilayer superconductors

### v2

@@ Abstract
We emphasize that the renormalized Josephson couplings in Eqs. (14) and (15) are derived in a mathematically rigorous manner, without reliance on model-specific parameters or physical assumptions, and are therefore generic. Then, the key point is whether Josephson phase fluctuations can be effectively excited. Typically, the Josephson coupling is weak, such that the gap of the Josephson phase mode $\omega_{L}(q=0)$ in Eq. (18)] lies below the SC gap, rendering it a thermally active low-energy excitation. Notably, an equally important factor is the phase stiffness $f_{s}$, which determines the available phase space for bosonic phase fluctuations entering the fluctuation integrals [Eq. (17)]. In bulk systems, the phase stiffness is generally large, strongly suppressing phase fluctuations. By contrast, in low-dimensional systems the phase stiffness is significantly reduced, owing both to a smaller Fermi velocity and to the intrinsically enhanced disorder associated with reduced dimensionality [71, 72]. Increasing disorder and reducing the Fermi velocity, as shown in Figs. 3(c) and (d), respectively, lead to a suppression of $T_{s}$, $T_{c}$, $T_{\eta}$, while markedly increasing the separation among these characteristic temperatures. As a result, the hierarchy of SC decoherence scales becomes increasingly pronounced. Among these effects, the suppression of gap closing temperature $T_{s}$ originates from the enhanced NG phase fluctuations (as discussed in Refs. [58, 59, 60]), whereas all the remaining Josephson features are primarily driven by the enhanced Josephson phase fluctuations (i.e., decoherence).
@@
Discussion.—Beyond the conventional single-energy-scale paradigm of the Josephson CPR, the present study demonstrates that phase decoherence constitutes an independent and more fragile energy scale in low-dimensional Josephson systems, and hence, Josephson physics itself in this case splits into multiple decoherence regimes, including diode nonreciprocity as the most fragile scale. We therefore predict that the forward and backward critical currents in low-dimensional Josephson-diode systems, $|I_{c}^{+}(T)|$ and $|I_{c}^{-}(T)|$, to merge as temperature increases, before vanishing altogether at $T_{c}$, as shown in Fig. 3(b). A similar tendency appears to be present in a wide range of recent experiments (e.g., Refs. [18, 19]), where the data exhibit a temperature evolution qualitatively consistent with that shown in Fig. 3(b).
@@
It should be emphasized that the vanishing of the Josephson critical current and the disappearance of the diode effect correspond to exponential thermal crossovers, rather than conventional thermodynamic phase transitions of first or second order. Both the diode efficiency $\eta(T)$ and the Josephson critical current $I_{c}(T)$ shown in Fig 2(b) exhibit behaviors with continuous temperature derivatives in the vicinity of their respective disappearance temperatures, in contrast to the mean-field scenario shown in Fig 2(a), where discontinuities in temperature derivatives are expected. Consequently, these crossovers do not exhibit thermodynamic signatures such as latent heat or a jump in the specific heat, in contrast to the SC gap-closing transition, which is a genuine second-order phase transition [45, 46]. Such critical behaviors of the Josephson properties therefore constitute a key distinguishing feature between mean-field and fluctuation-based descriptions of Josephson systems.
@@
The present work provides a tractable description of Josephson decoherence. Owing to the structural similarity [51, 60], this tractable framework can be naturally extended to layered superconductors with multilayer structures and Josephson coupling $J_{c}\cos(\theta_{n+1}-\theta_{n})$ between different layers (where $\theta_{n}$ is the SC phase in the $n$-th layer), such as cuprates [31, 32, 33, 34, 35, 36, 37, 38] and recently discovered nickelates [39, 40, 41, 42, 43, 44], particularly for their $c$-axis transport behavior. In this context, one can anticipate the disappearance of $c$-axis superconductivity (zero-resistance) while the SC gap persists above a characteristic critical temperature. Moreover, owing to the close analogy in the underlying derivation of the excitations, similar considerations also apply to SC transmon qubits [73, 74, 75]. Our results indicate that higher-order Josephson harmonics are increasingly fragile against phase decoherence. This effect is expected to be relevant for the stability/coherence of Josephson-based SC qubits, where such higher harmonics play an essential role in determining the qubit spectrum and anharmonicity.

### v3

@@ Abstract
We emphasize that the renormalized Josephson couplings in Eqs. (13) and (14) are derived in a mathematically rigorous manner, without reliance on model-specific parameters or physical assumptions, and are therefore generic. Then, the key point is whether Josephson phase fluctuations can be effectively excited. Typically, the Josephson coupling is weak, such that the gap of the Josephson phase mode $\omega_{L}(q=0)$ in Eq. (17)] lies below the SC gap, rendering it a thermally active low-energy excitation. Notably, an equally important factor is the phase stiffness $f_{s}$, which determines the available phase space for bosonic phase fluctuations entering the fluctuation integrals [Eq. (16)]. In bulk systems, the phase stiffness is generally large, strongly suppressing phase fluctuations. By contrast, in low-dimensional systems the phase stiffness is significantly reduced, owing both to a smaller Fermi velocity and to the intrinsically enhanced disorder associated with reduced dimensionality [57, 58]. Increasing disorder and reducing the Fermi velocity, as shown in Figs. 3(a) and (b), respectively, lead to a suppression of $T_{s}$, $T_{c}$, $T_{\eta}$, while markedly increasing the separation among these characteristic temperatures. As a result, the hierarchy of SC decoherence scales becomes increasingly pronounced. Among these effects, the suppression of gap closing temperature $T_{s}$ originates from the enhanced NG phase fluctuations (as discussed in Refs. [47, 48, 49]), whereas all the remaining Josephson features are primarily driven by the enhanced Josephson phase fluctuations (i.e., decoherence).
@@
Discussion.—Beyond the conventional single-energy-scale paradigm of the Josephson physics, the present study demonstrates that phase decoherence constitutes an independent and more fragile energy scale in low-dimensional Josephson systems, and hence, Josephson physics itself in this case splits into multiple decoherence regimes, including diode nonreciprocity as the most fragile scale. We therefore predict that the forward and backward critical currents in low-dimensional Josephson-diode systems, $|I_{c}^{+}(T)|$ and $|I_{c}^{-}(T)|$, to merge as temperature increases, before vanishing altogether at $T_{c}$, as shown in Fig. 3(c). A similar tendency appears to be present in a wide range of recent experiments (e.g., Refs. [24, 23, 53, 54, 55, 56]), where the data [see Fig. 3(d) and (e)] exhibit a temperature evolution qualitatively consistent with theoretical prediction.
@@
It should be emphasized that the vanishing of the Josephson critical current and the disappearance of the diode effect correspond to exponential thermal crossovers, rather than conventional thermodynamic phase transitions of first or second order. Both the diode efficiency $\eta(T)$ and the Josephson critical current $I_{c}(T)$ shown in Fig. 2(b) exhibit behaviors with continuous temperature derivatives in the vicinity of their respective disappearance temperatures, in contrast to the mean-field scenario shown in Fig. 2(a), where discontinuities in temperature derivatives are expected. Consequently, these crossovers do not exhibit thermodynamic signatures such as latent heat or a jump in the specific heat, in contrast to the SC gap-closing transition, which is a genuine second-order phase transition [33, 34]. Such critical behaviors of the Josephson properties thus constitute a key distinguishing feature between mean-field and fluctuation-based descriptions of Josephson systems. The proposed smooth-SC-decoherence mechanism follows analytically from the structure of both quantum and thermal phase fluctuations in low dimensions and does not rely on microscopic pairing details, symmetry-specific mechanisms, or fine tuning. It is fundamentally distinct from the BKT transition, the well-known mechanism for destroying SC phase coherence in two dimensions, which acts on the gapless global phase through vortex unbinding. By contrast, the Josephson phase mode is gapped due to interlayer phase locking and therefore does not participate in BKT physics.
@@
The developed theory provides a tractable description of SC decoherence processes. Owing to the structural similarity [37, 49], this framework can be potentially extended to layered superconductors with multilayer structures and Josephson coupling $J_{c}\cos(\theta_{n+1}-\theta_{n})$ between different layers (where $\theta_{n}$ is the SC phase in the $n$-th layer), such as cuprates [59, 60, 61, 62, 63, 64, 65, 66] and recently discovered nickelates [67, 68, 69, 70, 71, 72], particularly for their $c$-axis transport behavior. In this context, one can anticipate the disappearance of $c$-axis superconductivity (zero-resistance) while the SC gap persists above a characteristic critical temperature. Moreover, owing to the close analogy in the underlying derivation of the excitations, similar considerations also apply to SC transmon qubits [73, 74, 75]. Our results indicate that higher-order Josephson harmonics are increasingly fragile against decoherence. This effect is expected to be relevant for the stability/coherence of Josephson-based SC qubits, where such higher harmonics play an essential role in determining the qubit spectrum and anharmonicity.